OpenStudy (anonymous):

When solving this: f(x) = 2x-5 [f(x+h)-f(x)]/h Explain why I should get this: [2(x+h) -5(2x-5)]/h Rather than this: [(2x-5+h) -(2x-5)]/h I have seen the correct answer, but I want to know *WHY* it's supposed to turn out that way. If I am putting X+H in wherever I see X, shouldn't it be the latter? In [2(x+h) -5(2x-5)]/h, where is the (2x-5) after the -5 coming from?

OpenStudy (shamil98):

f(x) = 2x-5 [f(x+h)-f(x)]/h <--- this what goes into the value of x in the function above so it goes into the value of x, for example if f(x) = x +5 then it would be 2(x+5) - 5

OpenStudy (shamil98):

the -5 is coming from f(x) = 2x -5 let's say for x = (x + h) it would be 2(x+h) -5 do you understand it?..